# Roberto Dale,22footnotetext: Corresponding author.    Jesús M. Gandía    Juan Antonio Morales-Lladosa    and Ramon Lapiedra

### v1

@@ 3.3 Recovering the monopole and dipole
Table 1: Monopole and Dipole values used in the expected values calculations involved in Eqs. (2.1)–(2.3). As pointed out in [74], the uncertainty on the Planck dipole amplitude does not include the $0.02\%$ uncertainty of the temperature of the CMB monopole.
@@ 3.4 Selection criteria and sampling validation
The calculation of the aforementioned $F({R\boldsymbol{\hat{x}}})F({R\boldsymbol{\hat{y}}}),\alpha\right>$ quantities for a fixed value of $\alpha$ involves a certain number of angles $\alpha_{ij}$, with $6^{\circ}<\alpha_{ij}\leq 180^{\circ}$ included in a narrow interval, centred on $\alpha$, of width $\delta\alpha>0$, that is ${\alpha_{ij}}\in\left[{\alpha-\delta\alpha,\alpha+\delta\alpha}\right]$. Notice that $\alpha_{ij}>6^{\circ}$ is a ‘safety’ confidence margin to discard the causal connected directions at the last scattering surface. This width $\delta\alpha$ is related to the mean spacing between pixels, defined by the map resolution. For instance to build correlation functions, WMAP established for a $N_{\rm side}=16$ ($r=4$), which a mean spacing of $\thicksim 3^{\circ}$, a bin size of $6^{\circ}$. In that case, a total of 30 bins between $0^{\circ}$ and $180^{\circ}$ were settled [57]. However, since our computing capacity allows us to increase the resolution of the downgraded maps, we have degraded the original maps to a resolution of $r=6$ (49,152 pixels per map) instead of $r=4$ (3,072 pixels per map). Two short remarks: a resolution of $r=6$ is the resolution we can find in Planck correlation functions papers (see [58]) and taking into account that the mean spacing, in such a case, is $\thicksim 1^{\circ}$ (bin size $\thicksim 2^{\circ}$) –see Table 4 in Appendix A– a total of 90 bins were adopted for numerical calculations.
@@ 3.5 Expected values calculation
To get the $P\left({X,Y,\alpha}\right)$ probabilities appearing in expression (2.3) we will first define the probability $p_{\pm}(\boldsymbol{\hat{x}})$ of getting the dichotomised value ±$\pm 1$ of function $F({\boldsymbol{\hat{x}}})$ that will take into account the absolute error in $\delta T(\boldsymbol{\hat{x}})$ quantities. This absolute error, hereafter $\varepsilon_{a}(\delta T)$, is built up from the satellites data sets and their public information (see Subsection 3.6).
Henceforth, for a fixed direction $\boldsymbol{\hat{x}}$, the $\varepsilon_{a}(\delta T)$ will represent the absolute error in $\delta T$. Based on the $\delta T$ interval defined by $[\delta T-\varepsilon_{a}(\delta T),\delta T+\varepsilon_{a}(\delta T)]$ for a fixed direction $\boldsymbol{\hat{x}}$, the aforementioned $p_{\pm}(\boldsymbol{\hat{x}})$ are defined as follows: if we call $\varepsilon_{+}$ the positive section of the interval $[\delta T-\varepsilon_{a}(\delta T),\delta T+\varepsilon_{a}(\delta T)]$ the probability $p_{+}(\boldsymbol{\hat{x}})$ can be defined as
$p_{+}(\boldsymbol{\hat{x}})\equiv\frac{\varepsilon_{+}}{2\varepsilon_{a}(\delta T)},$ (3.1)
@@
$p_{-}(\boldsymbol{\hat{x}})\equiv\frac{{2\varepsilon_{a}(\delta T)}-\varepsilon_{+}}{2\varepsilon_{a}(\delta T)},$ (3.2)
@@
Now, based on these expressions, we can define, for a single vector pair $(\boldsymbol{\hat{x}},\boldsymbol{\hat{y}})$ the probabilities $p_{XY}(\boldsymbol{\hat{x}},\boldsymbol{\hat{y}})=p_{X}(\boldsymbol{\hat{x}})\cdot p_{Y}(\boldsymbol{\hat{y}})$, which represent the probabily of getting a XY value. In other words, the probability of obtaining each of the four possible values (+++++ +, ++-+ -, +-+- + and --- - ).
And finally, the probabilities $P(X,Y,\alpha)$ that appear in the equation (2.3), supposing that we have a total number of $N_{\alpha}$ pairs of vectors $(\boldsymbol{\hat{x}},\boldsymbol{\hat{y}})$ subtending certain angle $\alpha$, are
$P\left({X,Y,\alpha}\right)=\sum\limits_{i=1}^{{N_{\alpha}}}{\frac{{{p_{XY}(\boldsymbol{\hat{x}},\boldsymbol{\hat{y}})_{i}}}}{{{N_{\alpha}}}}},$ (3.3)
where $X=\pm 1$, $Y=\pm 1$. That completes the main ingredients for the calculations of the expected values (2.3). Nevertheless in the next section, a few details concerning the error calculation will be discussed.
@@ 3.6 Notes on error calculation
Once we have the error in each of the directions, it is immediate to obtain the absolute error in $\delta T(\boldsymbol{\hat{x}})$, that is the $\varepsilon_{a}(\delta T)$ quantities.
@@
${\varepsilon_{XY}}\left(\alpha\right)=Z\sqrt{\frac{{P\left({X,Y,\alpha}\right)\left[{1-P\left({X,Y,\alpha}\right)}\right]}}{{{N_{\alpha}}}}},$ (3.4)
where $Z$ stands for $z$-score whose chosen value was $Z=1.96$ at $2\sigma$ level. From these values we can calculate the absolute error in the expected values defined in (2.3), by means of
${\varepsilon_{a}}\left[\left\langle\alpha\right\rangle\right]={\left({{\varepsilon_{++}}^{2}+{\varepsilon_{+-}}^{2}+{\varepsilon_{-+}}^{2}+{\varepsilon_{--}}^{2}}\right)^{{\textstyle{1\over 2}}}}.$ (3.5)
The last step that remains for us, to complete the error calculation, is the one corresponding to the magnitude $\left|{\left\langle C\right\rangle}\right|$ detailed in the inequality (2.1). Proceeding in the same way as in our previous step, we have finally
@@ 4 Summary of computing tasks
In this section, we provide a concise review of the computational techniques employed in obtaining the results related to our evaluation of the CHSH inequalities (2.1). Before presenting the final outcomes, we will outline key details of the computational process, which was implemented in several stages.
@@
    A second code processes a total of seven downgraded maps, three from WMAP and four from Planck. This algorithm calculates the expected values, considering the elements we have described in the previous sections. In order to optimise the computation times, this code performs the operations based on a triangular matrix $152\times 49,$ 152 × 49 , 152 (total number of pixels for the resolution $r=6$) containing all the possible angles $\alpha$. Note that as the resolution increases, this technique based on the aforementioned triangular matrix is no longer viable, since for example, for a resolution $r=9$, such as the original WMAP resolution, more than $70,000$ GB of RAM would be necessary. An outrageous amount of memory, beyond our reach.
@@
    $\left|{\cos\alpha_{3}-\cos\beta\cos\alpha_{1}}\right|\leq{\sin\beta\sin\alpha_{1}},$ (4.1a)
@@
    $\left|{\cos\alpha_{4}-\cos\beta\cos\alpha_{2}}\right|\leq{\sin\beta\sin\alpha_{2}}.$ (4.1b)
@@ 5 NASA-WMAP summarised outcomes
5.1 The expected values $\alpha\right>$
In Figure 1, the expected value curves (2.3) derived from the analysis are presented in two panels, corresponding to the distinct downgrade methodologies employed. At first glance, two noteworthy properties can be emphasised: (i) the three curves, representing the Q, V, and W bands, are nearly identical, and (ii) consistent with results reported in analogous studies utilising data from the COBE satellite [10], the curves exhibit a pattern resembling a weighted cosine function, closely approximating the form $A\cos\left(\alpha\right)$. The most significant deviations from this form are observed for angles $\alpha\lesssim 18^{\circ}$ and $\alpha\gtrsim 165^{\circ}$ degrees. However, while the approximate amplitude reported in the COBE study was $0.60$, the amplitude observed in this analysis is notably higher, around $0.85$. This discrepancy motivated us to conduct a detailed investigation, thoroughly re-examining the potential violation of the (2.1) inequalities.
@@ 5.1 The expected values $\alpha\right>$
Downgrade Method Four-angle set Q Band V Band W Band Q, V and W
@@
Neighbour Average 133 49 137 139 2.4208 2.4222 2.4224 0.0009
133 49 135 141 2.4193 2.4207 2.4208 0.0009
129 45 137 139 2.4184 2.4197 2.4199 0.0009
131 45 133 141 2.4183 2.4196 2.4198 0.0009
133 51 137 141 2.4171 2.4184 2.4186 0.0009
Spherical Harmonics 133 49 137 139 2.4129 2.4146 2.4157 0.0009
133 49 135 141 2.4115 2.4132 2.4143 0.0009
129 45 137 139 2.4104 2.4122 2.4133 0.0009
131 45 133 141 2.4103 2.4121 2.4132 0.0009
133 51 137 141 2.4094 2.4111 2.4121 0.0009
Figure 2: As shown in Figure 1, the left and right panels present curves of the expected value $\alpha\right>$ calculated from Planck maps degraded using the average of neighbours method and spherical harmonic expansion, respectively. In both cases, four curves are plotted (one for each pipeline: COMMANDER, dash-dotted black line; NILC, dashed red line; SEVEM, dotted blue line; SMICA, solid green line), along with a fifth curve representing the $0.85\,\cos\left(\alpha\right)$ function (dashed gray line).
@@ 6 ESA-Planck summarised outcomes
6.1 The expected values $\alpha\right>$
In this case, Figure 2 displays the expected value curves $\alpha\right>$ calculated using the same two degradation methods previously applied to WMAP data, for each of the aforementioned processing pipelines. Several general characteristics can be observed in the figure: (i) unlike the curves in Figure 1, where it is difficult to visually discern which curve corresponds to each band, in this instance, the differences between the pipelines are clearly distinguishable, and (ii) as was the case with WMAP, the curves exhibit a good fit to the function $0.85\,\cos\left(\alpha\right)$. However, while, as with WMAP, noticeable discrepancies are observed between the four curves and the referenced function within a small range at the start and end of the curves, additional subtle but discernible differences are also present in other regions, depending on the specific curve under consideration.
@@ 6.2 Comparison of Expected Values from Different ESA-Planck Pipelines
$\Delta_{XY}\equiv 2\left|\frac{{\left\langle\alpha\right\rangle_{X}-\left\langle\alpha\right\rangle_{Y}}}{{\left\langle\alpha\right\rangle_{X}+\left\langle\alpha\right\rangle_{Y}}}\right|,$ (6.1)
@@
COMMANDER 43 135 45 47 2,1949 0,0009 43 135 45 47 2.1956 0.0009
133 49 135 141 2.1932 0.0009 133 49 135 141 2.1938 0.0009
133 49 137 139 2.1929 0.0009 133 49 137 139 2.1935 0.0009
43 137 45 49 2.1929 0.0009 131 45 133 141 2.1935 0.0009
131 45 133 141 2.1929 0.0009 43 137 45 49 2.1935 0.0009
NILC 43 135 45 47 2.4353 0.0010 43 135 45 47 2.4356 0.0010
133 49 137 139 2.4331 0.0009 133 49 137 139 2.4333 0.0009
43 137 45 49 2.4324 0.0010 43 137 45 49 2.4327 0.0010
133 49 135 141 2.4317 0.0009 133 49 135 141 2.4320 0.0009
43 139 47 49 2.4308 0.0010 43 139 47 49 2.4311 0.0010
SEVEM 43 135 45 47 2.3216 0.0010 43 135 45 47 2.3226 0.0010
133 49 137 139 2.3193 0.0009 133 49 137 139 2.3203 0.0009
43 137 45 49 2.3190 0.0010 43 137 45 49 2.3199 0.0010
133 49 135 141 2.3189 0.0009 133 49 135 141 2.3199 0.0009
131 45 133 141 2.3184 0.0009 131 45 133 141 2.3193 0.0009
SMICA 43 135 45 47 2.2747 0.0010 43 135 45 47 2.2756 0.0010
43 137 45 49 2.2723 0.0010 43 137 45 49 2.2732 0.0010
133 49 137 139 2.2716 0.0009 133 49 137 139 2.2725 0.0009
133 49 135 141 2.2711 0.0009 133 49 135 141 2.2721 0.0009
131 39 133 135 2.2708 0.0009 131 39 133 135 2.2718 0.0009
@@
    For a given pipeline, both map degradation methods yield the same five best tetrads. The order of these tetrads is also largely preserved, with the exception of the fourth and fifth positions, which are reversed in the case of COMMANDER.
@@
    In contrast to the data from WMAP, where the tetrads were identical across all three frequency bands, small differences are observed in both the values of the angles and their ordering. For example, the tetrad $\left\{131^{\circ},39^{\circ},133^{\circ},135^{\circ}\right\}$ is unique to the SMICA pipeline. Nevertheless, the tetrad $\left\{43^{\circ},135^{\circ},45^{\circ},47^{\circ}\right\}$ consistently yields the highest value across all cases.
@@
    As can be inferred from the curves in Figure 3, the largest discrepancies occur between the COMMANDER and NILC pipelines, with a difference of approximately $10\%$. The SEVEM and SMICA pipelines yield the closest results, differing by approximately $2\%$.
  5. (e)

### v2

@@ 3.3 Recovering the monopole and dipole
Table 1: Monopole and Dipole values used in the expected values calculations involved in Eqs. (1)–(3). As pointed out in [74], the uncertainty on the Planck dipole amplitude does not include the $0.02\%$ uncertainty of the temperature of the CMB monopole.
@@ 3.4 Selection criteria and sampling validation
The calculation of the aforementioned $\left<F({R\boldsymbol{\hat{x}}})F({R\boldsymbol{\hat{y}}}),\alpha\right$> quantities for a fixed value of $\alpha$ involves a certain number of angles $\alpha_{ij}$, with $6^{\circ}<\alpha_{ij}\leq 180^{\circ}$ included in a narrow interval, centred on $\alpha$, of width $\delta\alpha>0$, that is ${\alpha_{ij}}\in\left[{\alpha-\delta\alpha,\alpha+\delta\alpha}\right]$. Notice that $\alpha_{ij}>6^{\circ}$ is a ‘safety’ confidence margin to discard the causal connected directions at the last scattering surface. This width $\delta\alpha$ is related to the mean spacing between pixels, defined by the map resolution. For instance to build correlation functions, WMAP established for a $N_{\rm side}=16$ ($r=4$), which a mean spacing of $\thicksim 3^{\circ}$, a bin size of $6^{\circ}$. In that case, a total of 30 bins between $0^{\circ}$ and $180^{\circ}$ were settled [57]. However, since our computing capacity allows us to increase the resolution of the downgraded maps, we have degraded the original maps to a resolution of $r=6$ (49,152 pixels per map) instead of $r=4$ (3,072 pixels per map). Two short remarks: a resolution of $r=6$ is the resolution we can find in Planck correlation functions papers (see [58]) and taking into account that the mean spacing, in such a case, is $\thicksim 1^{\circ}$ (bin size $\thicksim 2^{\circ}$) –see Table 4 in Appendix A– a total of 90 bins were adopted for numerical calculations.
@@ 3.5 Expected values calculation
[…]
$p_{+}(\boldsymbol{\hat{x}})\equiv\frac{\varepsilon_{+}}{2\varepsilon_{a}(\delta T)},$ (4)
@@
$p_{-}(\boldsymbol{\hat{x}})\equiv\frac{{2\varepsilon_{a}(\delta T)}-\varepsilon_{+}}{2\varepsilon_{a}(\delta T)},$ (5)
@@
[…]
$P\left({X,Y,\alpha}\right)=\sum\limits_{i=1}^{{N_{\alpha}}}{\frac{{{p_{XY}(\boldsymbol{\hat{x}},\boldsymbol{\hat{y}})_{i}}}}{{{N_{\alpha}}}}},$ (6)
where $X=\pm 1$, $Y=\pm 1$. That completes the main ingredients for the calculations of the expected values (3). Nevertheless in the next section, a few details concerning the error calculation will be discussed.
@@ 3.6 Notes on error calculation
[…]
@@
${\varepsilon_{XY}}\left(\alpha\right)=Z\sqrt{\frac{{P\left({X,Y,\alpha}\right)\left[{1-P\left({X,Y,\alpha}\right)}\right]}}{{{N_{\alpha}}}}},$ (7)
where $Z$ stands for $z$-score whose chosen value was $Z=1.96$ at $2\sigma$ level. From these values we can calculate the absolute error in the expected values defined in (3), by means of
${\varepsilon_{a}}\left[\left\langle\alpha\right\rangle\right]={\left({{\varepsilon_{++}}^{2}+{\varepsilon_{+-}}^{2}+{\varepsilon_{-+}}^{2}+{\varepsilon_{--}}^{2}}\right)^{{\textstyle{1\over 2}}}}.$ (8)
The last step that remains for us, to complete the error calculation, is the one corresponding to the magnitude $\left|{\left\langle C\right\rangle}\right|$ detailed in the inequality (1). Proceeding in the same way as in our previous step, we have finally
$\left.\begin{aligned} {\varepsilon_{a}}\left({\left|{\left\langle C\right\rangle}\right|}\right)&=\left[{\varepsilon_{a}}^{2}\left({\left\langle{{\alpha_{1}}}\right\rangle}\right)+{\varepsilon_{a}}^{2}\left({\left\langle{{\alpha_{2}}}\right\rangle}\right)+{\varepsilon_{a}}^{2}\left({\left\langle{{\alpha_{3}}}\right\rangle}\right)\right.+\left.{\varepsilon_{a}}^{2}\left({\left\langle{{\alpha_{4}}}\right\rangle}\right)\right]^{{\textstyle{1\over 2}}}.\end{aligned}\right$. (9)
@@ 4 Summary of computing tasks
In this section, we provide a concise review of the computational techniques employed in obtaining the results related to our evaluation of the CHSH inequalities (1). Before presenting the final outcomes, we will outline key details of the computational process, which was implemented in several stages.
@@
    A second code processes a total of seven downgraded maps, three from WMAP and four from Planck. This algorithm calculates the expected values, considering the elements we have described in the previous sections. In order to optimise the computation times, this code performs the operations based on a triangular matrix $49,152\times 49,152$ (total number of pixels for the resolution $r=6$) containing all the possible angles $\alpha$. Note that as the resolution increases, this technique based on the aforementioned triangular matrix is no longer viable, since for example, for a resolution $r=9$, such as the original WMAP resolution, more than $70,000$ GB of RAM would be necessary. An outrageous amount of memory, beyond our reach.
@@
    $\left|{\cos\alpha_{3}-\cos\beta\cos\alpha_{1}}\right|\leq{\sin\beta\sin\alpha_{1}},$ (10a)
@@
    $\left|{\cos\alpha_{4}-\cos\beta\cos\alpha_{2}}\right|\leq{\sin\beta\sin\alpha_{2}}.$ (10b)
@@ 5 NASA-WMAP summarised outcomes
5.1 The expected values $\left<\alpha\right$>
In Figure 1, the expected value curves (3) derived from the analysis are presented in two panels, corresponding to the distinct downgrade methodologies employed. At first glance, two noteworthy properties can be emphasised: (i) the three curves, representing the Q, V, and W bands, are nearly identical, and (ii) consistent with results reported in analogous studies utilising data from the COBE satellite [10], the curves exhibit a pattern resembling a weighted cosine function, closely approximating the form $A\cos\left(\alpha\right)$. The most significant deviations from this form are observed for angles $\alpha\lesssim 18^{\circ}$ and $\alpha\gtrsim 165^{\circ}$ degrees. However, while the approximate amplitude reported in the COBE study was $0.60$, the amplitude observed in this analysis is notably higher, around $0.85$. This discrepancy motivated us to conduct a detailed investigation, thoroughly re-examining the potential violation of the (1) inequalities.
@@ 5.1 The expected values $\left<\alpha\right$>
Downgrade
Method
Four-angle set Q Band V Band W Band Q, V and W
@@
Neighbour
Average
133 135 137 45 2.4248 2.4262 2.4263 0.0009
133 135 139 47 2.4234 2.4248 2.4249 0.0009
133 137 139 49 2.4208 2.4222 2.4224 0.0009
133 135 141 49 2.4193 2.4207 2.4208 0.0009
131 135 141 47 2.4188 2.4202 2.4203 0.0009
Spherical
Harmonics
133 135 137 45 2.4165 2.4183 2.4194 0.0009
133 135 139 47 2.4153 2.4171 2.4182 0.0009
133 137 139 49 2.4129 2.4146 2.4157 0.0009
133 135 141 49 2.4115 2.4132 2.4143 0.0009
131 135 141 47 2.4108 2.4126 2.4137 0.0009
Figure 2: As shown in Figure 1, the left and right panels present curves of the expected value $\left<\alpha\right$> calculated from Planck maps degraded using the average of neighbours method and spherical harmonic expansion, respectively. In both cases, four curves are plotted (one for each pipeline: COMMANDER, dash-dotted black line; NILC, dashed red line; SEVEM, dotted blue line; SMICA, solid green line), along with a fifth curve representing the $0.85\,\cos\left(\alpha\right)$ function (dashed gray line).
@@ 6 ESA-Planck summarised outcomes
6.1 The expected values $\left<\alpha\right$>
In this case, Figure 2 displays the expected value curves $\left<\alpha\right$> calculated using the same two degradation methods previously applied to WMAP data, for each of the aforementioned processing pipelines. Several general characteristics can be observed in the figure: (i) unlike the curves in Figure 1, where it is difficult to visually discern which curve corresponds to each band, in this instance, the differences between the pipelines are clearly distinguishable, and (ii) as was the case with WMAP, the curves exhibit a good fit to the function $0.85\,\cos\left(\alpha\right)$. However, while, as with WMAP, noticeable discrepancies are observed between the four curves and the referenced function within a small range at the start and end of the curves, additional subtle but discernible differences are also present in other regions, depending on the specific curve under consideration.
@@ 6.2 Comparison of Expected Values from Different ESA-Planck Pipelines
$\Delta_{XY}\equiv 2\left|\frac{{\left\langle\alpha\right\rangle_{X}-\left\langle\alpha\right\rangle_{Y}}}{{\left\langle\alpha\right\rangle_{X}+\left\langle\alpha\right\rangle_{Y}}}\right|,$ (11)
@@
COMMANDER 133 135 137 45 2.1960 0,0009 133 135 137 45 2.1966 0.0009
133 135 139 47 2.1950 0.0009 133 135 139 47 2.1956 0.0009
43 45 47 135 2.1949 0.0009 43 45 47 135 2.1956 0.0009
41 45 47 133 2.1937 0.0009 41 45 47 133 2.1943 0.0009
133 135 141 49 2.1932 0.0009 133 135 141 49 2.1938 0.0009
NILC 133 135 137 45 2.4376 0.0010 133 135 137 45 2.4378 0.0009
133 135 139 47 2.4360 0.0009 133 135 139 47 2.4362 0.0009
43 45 47 135 2.4353 0.0010 43 45 47 135 2.4356 0.0010
41 45 47 133 2.4332 0.0010 41 45 47 133 2.4334 0.0010
133 137 139 49 2.4331 0.0009 133 137 139 49 2.4333 0.0010
SEVEM 133 135 137 45 2.3233 0.0009 133 135 137 45 2.3243 0.0009
133 135 139 47 2.3220 0.0009 133 135 139 47 2.3229 0.0009
43 45 47 135 2.3216 0.0010 43 45 47 135 2.3226 0.0010
41 45 47 133 2.3201 0.0010 41 45 47 133 2.3211 0.0010
133 137 139 49 2.3193 0.0009 133 137 139 49 2.3203 0.0009
SMICA 133 135 137 45 2.2755 0.0009 133 135 137 45 2.2764 0.0009
43 45 47 135 2.2747 0.0010 43 45 47 135 2.2756 0.0010
133 135 139 47 2.2740 0.0009 133 135 139 47 2.2749 0.0009
41 45 47 133 2.2730 0.0010 41 45 47 133 2.2740 0.0010
133 137 139 49 2.2716 0.0009 133 137 139 49 2.2725 0.0009
@@
    For all pipelines considered, both map degradation methods produce the same five best tetrads. Moreover, the ordering of these tetrads is largely preserved across the two degradation approaches.
@@
    In contrast to the WMAP data, where the tetrads were identical across all three frequency bands, small discrepancies in the ordering of angular values are observed in the current analysis. For instance, the tetrads $\left\{43^{\circ},45^{\circ},47^{\circ},135^{\circ}\right\}$ and $\left\{133^{\circ},135^{\circ},139^{\circ},47^{\circ}\right\}$ appear in reverse order in the SMICA pipeline. Nevertheless, the tetrad $\left\{133^{\circ},135^{\circ},137^{\circ},45^{\circ}\right\}$ consistently yields the highest value across all cases.